# Why Transformers Need Adam: A Hessian Perspective

### v1

@@ C.1 Implementation Details on SLQ
We provide a simple PyTorch implementation of SLQ. The only query SLQ makes to the neural network is the Hessian vector product, which is attained using the auto-differentiation framework [Pearlmutter, 1994]. To assure the accuracy of the Lanczos algorithm, we remove all the randomness in the forward and backward passes, including: data shuffling order, data augmentation, and dropout, etc.. Since Flash Attention [Dao et al., 2022] does not support the calculation of Hessian-vector product, we implement all attention blocks in the naive way. For the calculation of the blockwise Hessian spectrum $\nabla^{2}\mathcal{L}(w_{l})$, we sample $u_{l}\sim N(0,I_{d_{l}\times d_{l}})$ and set $v_{l}=u_{l}/\|u_{l}\|_{2}\in\mathbb{R}^{d_{l}}$. Then we run Algorithm 5 by taking $\nabla^{2}\mathcal{L}(w_{l})$ and $v_{l}$ as inputs. We choose the hyperparameters as $m=100$ and $n_{v}=10$ in all experiments. $\sigma$ is tuned based on visual effects. These hyperparameters are reported to reach highly accurate estimation with error $<10^{-14}$ [Ghorbani et al., 2019].
We now briefly discuss the computational cost of SLQ. The major computational expense of SLQ is the repeated Hessian-vector product operations in Lanczos algorithm in Step 3. Recall $\nabla^{2}\mathcal{L}(w)d=\frac{1}{n}\sum_{i=1}^{n}\nabla^{2}\mathcal{L}_{i}(w)d$, so each Hessian-vector product operation requires (i) calculating $\nabla^{2}\mathcal{L}_{i}(w)d$; (ii) repeating (i) on all data. We point out that (i) can be computed efficiently and precisely with just two backpropagation passes [Pearlmutter, 1994]. The major computational bottleneck lies in (ii) due to the large $n$. Our largest-scale experiment for Hesian spectrum is GPT2 (125M) on Openwebtext, where the number of tokens $n=9$ Billon. To calculate $\nabla^{2}\mathcal{L}(w)d$ on all these 9B tokens, it requires about 9 GPU days on eight A100-80GB GPUs. Since SLQ requires at least 1,000 times query of $\nabla^{2}\mathcal{L}(w)d$, a complete run of SLQ would take at least 9,000 days on eight A100-80GB GPUs, which is unaffordable. In this work, we use the largest possible batch size (with gradient accumulation tricks) to approximate $\nabla^{2}\mathcal{L}(w)$ under the constraints of GPU bandwidth and time limit. More detailed setup of SLQ are shown as follows.
  • ResNet18 (20M) and VGG16 (138M) on ImageNet. We use the code base of PyTorch Examples  [^4] https://github.com/pytorch/examples/blob/main/imagenet/main.py. We use batch size $=1024$. For the calculation of the blockwise Hessian spectra, we apply SLQ to all parameter blocks except for the BatchNorm layers. In total, it takes about 3 days on one V100 GPU to estimate all the blockwise Hessian spectra and the full Hessian spectrum.
@@
  • GPT2 (125M) on Openwebtext. We use the code base of NanoGPT. We use batch size $=245,760$ tokens. Due to the large number of parameters, we are not able to calculate the blockwise Hessian spectra for all parameter blocks. Instead, we apply SLQ to: the embedding layer; the output layer; the $1$-st, $4$-th, $8$-th, $12$-th attention blocks; and the $1$-st, $4$-th, $8$-th, $12$-th MLP blocks (note that the $12$-th attention and MLP blocks are the final ones). In total, it takes about 7 days on one A100 GPU to estimate all the blockwise Hessian spectra and the full Hessian spectrum.
We train all the models under the default configurations in all cases. For CNN training, we follow the rule of thumb to choose the learning rate of AdamW: we set the learning rate of AdamW to be 100×100$\times$100 × smaller than the recommended learning rate of SGD.
C.2 Implementation Details on Figure LABEL:fig_block_diagnal
@@
We use the code base of NanoGPT to train a decoder-only Transformers on 4 consecutive tokens randomly selected from Openwebtext. We set the model configuration as: context window = 2, number of heads = 2 and the embedding dimension = 4. In MLP layers, all widths equal to 4. We change the number of layers from 1 to 4. We remove all the Layer Norms in the model. The rest of the model configurations are set to their default values in the code base. We compute the Hessian on all the parameters blocks in attention and MLP layers. In Figure LABEL:fig_block_diagnal (a), the variables in three blue boxes corresponds to the parameters in Query and Key; Value and Projection; and MLP, respectively. Similarly for the rest of the figures.
@@
Due to the intensive overhead of computing and storing the whole Hessian, we have yet to check the block-diagonal structure on larger models. Rigorously speaking, so far we have not gotten sufficient evidence to claim this structure commonly holds in larger Transformers. It requires new numerical methods to efficiently check the block-diagonal Hessian structure in larger models without explicitly calculating them. We leave it as an interesting future direction.
For completeness, we remark that Collobert [2004], Roux et al. [2007], Martens and Grosse [2015] also observed the block-diagonal structure in (approximated) Hessian of small-scaled MLPs. We restate their findings in Figure LABEL:fig_block_diagnal_existing. [Collobert, 2004, Section 7] further theoretically proved that the block diagonal structure stems from (i) the layer-by-layer structure and (ii) the Cross-Entropy loss. These results suggest that block-diagonal structure might be common in NNs.
C.3 Implementation Details on the MLP experiments in Figure LABEL:fig_mlp_gap
@@
We train a 4-layer MLP on MNIST. We use batch size = 128 and width = 300, 128, and 64 for the hidden layers. We use ReLU activation. We change the degree of heterogeneity by scaling the output of each layer with constant $c\in\mathbb{N}$. We scale $c$ from 1 to 15. We record the average JS distance between the initial Hessian spectra of each pair of layers and find that the average JS distance increases as $c$ increases. As such, we define $c$ as the degree of block heterogeneity. For each $c$, we train SGD and Adam with default hyperparameters by grid-searching the learning rate from 1e-4 to 1e-1 and report the best test accuracy after 1 epoch.
$c$ 1 2 3 4 5 6 7 8 9-15
Blockwise distance 1e-3 1.61 33.99 65.32 260.77 298.39 351.69 353.48 $\operatorname{NaN}$
Table 3: Consider a 4-layer MLP on MNIST. $c$ denotes the scaler that is multiplied to the output of each layer. Blockwise distance denotes the average JS distance between the initial Hessian spectra of each pair of layers. We find that the average JS distance increases as $c$ increases. For $c\geq 9$, the calculation of JS distance encounters numerical instability due to the division of near-zero values and returns $\operatorname{NaN}$.
Appendix D More Experimental Results
D.1 Performance comparison of AdamW and SGD on Different Architectures
Here, we show the performance comparison of AdamW and SGD on different models.
@@ E.2 Proof of Theorem 1
Without loss of generality, we assume $h=0$. This is because minimizing $\mathcal{L}(w)=\frac{1}{2}w^{T}Hw-h^{T}w$ is equivalent to minimizing $\mathcal{L}(w)=\frac{1}{2}\left(w-w^{*}\right)^{T}H\left(w-w^{*}\right)$ where $w^{*}=H^{-1}h$. By a linear transformation $z=w-w^{*}$, BGD for minimizing $\frac{1}{2}\left(w-w^{*}\right)^{T}H\left(w-w^{*}\right)$ starting from $w^{0}$ is equivalent to BGD for minimizing $\frac{1}{2}z^{T}Hz$ starting from $z^{0}=w^{0}-w^{*}$. Thus we can assume $w^{*}=0$, or equivalently, $h=0$. When $h=0$, the update form of BGD becomes
$w^{t+1}=w^{t}-\eta D_{BGD}^{-1}Hw^{t},$
where $D_{BGD}=\operatorname{diag}\left(\left(\underbrace{\eta_{1},\cdots,\eta_{1}}_{d_{1}\text{times}},\cdots,\underbrace{\eta_{L},\cdots,\eta_{L}}_{d_{L}\text{times}}\right)^{T}\right)\in\mathbb{R}^{d\times d}.$
In the following, we will first prove the rate of BGD for each block and then combine them together to get the total rate. For the $l$-th block, we choose $\eta_{l}=\frac{1}{\lambda_{l,1}}$. The update rule of BGD for the $l$-th block becomes:
$w_{l}^{t+1}=w_{l}^{t}-\frac{1}{\lambda_{l,1}}H_{l}w_{l}^{t}$
Now we apply the standard convergence proof of GD to the $l$-th block. By descent lemma, we have
$[\mathcal{L}(w^{t+1})]_{l}$$\leq$ $[\mathcal{L}(w^{t})]_{l}+\langle w_{l}^{t+1}-w_{l}^{t},[\nabla\mathcal{L}(w^{t})]_{l}\rangle+\frac{\lambda_{l,1}}{2}\|w_{l}^{t+1}-w_{l}^{t}\|_{2}^{2}$ (11)
$\leq$ $[\mathcal{L}(w^{t})]_{l}+\langle w_{l}^{t+1}-w_{l}^{t},H_{l}w_{l}^{t}\rangle+\frac{\lambda_{l,1}}{2}\eta^{2}w_{l}^{t}H_{l}^{T}H_{l}w_{l}^{t}$ (12)
$=$ $[\mathcal{L}(w^{t})]_{l}-\frac{1}{2\lambda_{l,1}}w_{l}^{t}H_{l}^{T}H_{l}w_{l}^{t}$ (13)
$\leq$ $[\mathcal{L}(w^{t})]_{l}-\frac{\lambda_{l,d_{l}}}{2\lambda_{l,1}}w_{l}^{t}H_{l}w_{l}^{t}$ (14)
$=$ $\left(1-\frac{1}{\kappa_{l}}\right)[\mathcal{L}(w^{t})]_{l}$ (15)
Substract both sides by $\mathcal{L}^{*}$ and we get $[\mathcal{L}(w^{t+1})]_{l}-[\mathcal{L}^{*}]_{l}\leq\left(1-\frac{1}{\kappa_{l}}\right)\left([\mathcal{L}(w^{t})]_{l}-[\mathcal{L}^{*}]_{l}\right)$. Summing up both sides over $l\in[L]$,we have
$\mathcal{L}(w^{t+1})-\mathcal{L}^{*}$ $=$ $\sum_{l=1}^{L}[\left(\mathcal{L}(w^{t+1})]_{l}-[\mathcal{L}^{*}]_{l}\right)$
$\leq$ $\sum_{l=1}^{L}\left(1-\frac{1}{\kappa_{l}}\right)\left([\mathcal{L}(w^{t})]_{l}-[\mathcal{L}^{*}]_{l}\right)$
$\leq$ $\max_{l\in[L]}\left(1-\frac{1}{\kappa_{l}}\right)\sum_{l=1}^{L}\left([\mathcal{L}(w^{t})]_{l}-[\mathcal{L}^{*}]_{l}\right)$
$=$ $\max_{l\in[L]}\left(1-\frac{1}{\kappa_{l}}\right)\left(\mathcal{L}(w^{t})-\mathcal{L}^{*}\right).$
This concludes the proof of Theorem 1.
E.3 Proof of Theorem 2
Similarly as the proof of Theorem 1, we assume $h=0$. The update rule of Adam becomes
@@ E.3 Proof of Theorem 2
where $D_{Adam}^{0}=\operatorname{diag}(\nabla\mathcal{L}(w^{0})\circ\nabla\mathcal{L}(w^{0}))^{\frac{1}{2}}=\operatorname{diag}(|Hw^{0}|)$. We denote $d_{t}=\eta(D_{Adam}^{0})^{-1}Hw^{t}$ and thus we have $w^{t}=\frac{1}{\eta}H^{-1}D_{Adam}^{0}d^{t}$ and $w^{t+1}=w^{t}-d_{t}$. These relations also hold for each block by changing the notation to $H_{l}$ $w_{l}^{t}$, $D_{Adam}^{0}$, and $d_{l}^{t},$ etc.. Following the framework in [Sun and Ye, 2021], we try to bound the error yet to be optimized (a.k.a., cost-to-go) and the per-step improvement, respectively. The ratio of these two terms characterizes the rate of convergence. We now express both terms using $d_{l}^{t}$. For the cost-to-go term for the $l$-th block, we have
@@
$\leq$ $\left(\max_{l\in[L]}\frac{C_{l,2}^{2}}{C_{l,1}^{2}}\right)\kappa_{l},$ (20)
@@
$\leq$ $\sum_{l=1}^{L}\left(1-\frac{1}{\left(\max_{l\in[L]}\frac{C_{l,2}^{2}}{C_{l,1}^{2}}\right)\kappa_{l}}\right)\left([\mathcal{L}(w^{t})]_{l}-[\mathcal{L}^{*}]_{l}\right)$
$\leq$ $\max_{l\in[L]}\left(1-\frac{1}{\left(\max_{l\in[L]}\frac{C_{l,2}^{2}}{C_{l,1}^{2}}\right)\kappa_{l}}\right)\sum_{l=1}^{L}\left([\mathcal{L}(w^{t})]_{l}-[\mathcal{L}^{*}]_{l}\right)$
$=$ $\max_{l\in[L]}\left(1-\frac{1}{\left(\max_{l\in[L]}\frac{C_{l,2}^{2}}{C_{l,1}^{2}}\right)\kappa_{l}}\right)\left(\mathcal{L}(w^{t})-\mathcal{L}^{*}\right).$

### v2

@@ D.1 Implementation Details on SLQ and Training Configurations
We provide a simple PyTorch implementation of SLQ. The only query SLQ makes to the neural network is the Hessian vector product, which is attained using the auto-differentiation framework [70]. To assure the accuracy of the Lanczos algorithm, we remove all the randomness in the forward and backward passes, including: data shuffling order, data augmentation, and dropout, etc.. Since Flash Attention [22] does not support the calculation of Hessian-vector product, we implement all attention blocks in the naive way. For the calculation of the blockwise Hessian spectrum $\nabla^{2}\mathcal{L}(w_{l})$, we sample $u_{l}\sim N(0,I_{d_{l}\times d_{l}})$ and set $v_{l}=u_{l}/\|u_{l}\|_{2}\in\mathbb{R}^{d_{l}}$. Then we run Algorithm 5 by taking $\nabla^{2}\mathcal{L}(w_{l})$ and $v_{l}$ as inputs. We choose the hyperparameters as $m=100$ and $n_{v}=10$ in all experiments. $\sigma$ is tuned based on visual effects. These hyperparameters are reported to reach highly accurate estimation with error $<10^{-14}$ [32].
We now briefly discuss the computational cost of SLQ. The major computational expense of SLQ is the repeated Hessian-vector product operations in Lanczos algorithm in Step 3. Recall $\nabla^{2}\mathcal{L}(w)d=\frac{1}{n}\sum_{i=1}^{n}\nabla^{2}\mathcal{L}_{i}(w)d$, so each Hessian-vector product operation requires (i) calculating $\nabla^{2}\mathcal{L}_{i}(w)d$; (ii) repeating (i) on all data. We point out that (i) can be computed efficiently and precisely with just two backpropagation passes [70]. The major computational bottleneck lies in (ii) due to the large $n$. Our largest-scale experiment for Hesian spectrum is GPT2 (125M) on Openwebtext, where the number of tokens $n=9$ Billon. To calculate $\nabla^{2}\mathcal{L}(w)d$ on all these 9B tokens, it requires about 9 GPU days on eight A100-80GB GPUs. Since SLQ requires at least 1,000 times query of $\nabla^{2}\mathcal{L}(w)d$, a complete run of SLQ would take at least 9,000 days on eight A100-80GB GPUs, which is unaffordable. In this work, we use the largest possible batch size (with gradient accumulation tricks) to approximate $\nabla^{2}\mathcal{L}(w)$ under the constraints of GPU bandwidth and time limit. More detailed setup of SLQ are shown as follows.
  • ResNet18 (18M) and VGG16 (138M) on ImageNet. We use the code base of PyTorch Examples  [^5] https://github.com/pytorch/examples/blob/main/imagenet/main.py. We use batch size $=1024$. For the calculation of the blockwise Hessian spectra, we apply SLQ to all parameter blocks except for the BatchNorm layers. In total, it takes about 3 days on one V100 GPU to estimate all the blockwise Hessian spectra and the full Hessian spectrum.
@@
  • GPT2 (125M) on Openwebtext  [^9] https://huggingface.co/datasets/Skylion007/openwebtext. We use the code base of NanoGPT. We use batch size $=245,760$ tokens. Due to the large number of parameters, we are not able to calculate the blockwise Hessian spectra for all parameter blocks. Instead, we apply SLQ to: the embedding layer; the output layer; the $1$-st, $4$-th, $8$-th, $12$-th attention blocks; and the $1$-st, $4$-th, $8$-th, $12$-th MLP blocks (note that the $12$-th attention and MLP blocks are the final ones). In total, it takes about 7 days on one A100 GPU to estimate all the blockwise Hessian spectra and the full Hessian spectrum.
Training configuration.
In all cases, we train all the models under the default configurations in the above codebase. We grid-search the learning rates for SGD and Adam under the same budget and report the best result for each optimizer. We use the cosine-decay learning rate schedule for vision tasks. For SFT task, we use nanoGPT codebase. We first pre-train GPT2 on OpenwebText for 25B tokens and then fine-tune it on a subset of Alpaca Eval  [^10] https://huggingface.co/datasets/tatsu-lab/alpaca_eval.
D.2 Implementation Details on Figure LABEL:fig_block_diagnal
@@
We use the code base of NanoGPT to train decoder-only Transformers on 4 consecutive tokens randomly selected from Openwebtext. We set the model configuration as: context window = 2, number of heads = 2 and the embedding dimension = 4. In MLP layers, all widths equal to 4. We choose the number of layers to be 2, 4, and 8. We remove all the LayerNorms in the model. The rest of the model configurations are set to their default values in the code base. We compute the Hessian on all the parameters blocks in attention and MLP layers. In Figure LABEL:fig_block_diagnal (a), the variables each block corresponds to the parameters in Query and Key; Value and Projection; and MLP, respectively. For better visualization, We take the absolute value of each entry to distinguish non-zero values (including negatives) from those near 0 and then report the average value in each block. Similarly for the rest of the figures.
@@
Due to the intensive overhead of computing and storing the whole Hessian, we have yet to check the block-diagonal structure on larger models. Rigorously speaking, so far we have not gotten sufficient evidence to claim this structure commonly holds in larger Transformers. It requires new numerical methods to efficiently check the block-diagonal Hessian structure without explicitly calculating them. We leave it as an interesting future direction.
D.3 More Details on the MLP experiments in Figure LABEL:fig_mlp_gap
@@
We train a 4-layer MLP on MNIST. We use batch size = 128 and width = 300, 128, and 64 for the hidden layers. We use ReLU activation. We change the degree of heterogeneity by scaling the output of each layer with constant $c\in\mathbb{N}$. We scale $c$ from 1 to 15. For each $c$, we train SGD and Adam with default hyperparameters by grid-searching the learning rate from 1e-4 to 1e-1 and report the best test accuracy after 1 epoch.
@@ E.2 Proof of Theorem 1
We first show that $C_{l,2}$ and $C_{l,1}$ are non-zero w.p.1. under the random initialization in Assumption 1. We define set $S_{i}=\{w;h_{i}^{T}w=0\}$ where $h_{i}\in\mathbb{R}^{d}$ is the $i$-th row of $H$. Since $H$ is positive definite, there is at least one non-zero entry in $h_{i}$, $i\in[d]$. As such, $S_{i}$ is a $(d-1)$-dimensional subspace of $\mathbb{R}^{d}$ and thus has zero Lebesgue measure in $\mathbb{R}^{d}$. Since $w^{0}$ follows continuous distribution, we have $\Pr\left(\{w^{0};h_{i}^{T}w^{0}=0\}\right)=0$, for $i=[d]$. Then we have
$\Pr\left(\nabla\mathcal{L}(w^{0})\text{ has at least one zero entry}\right)$ $=$ $\Pr\left(Hw^{0}\text{ has at least one zero entry}\right)$ (11)
$=$ $\Pr\left(\cup_{i=1}^{d}\{w^{0};h_{i}^{T}w^{0}=0\}\right)$ (12)
$\leq$ $\sum_{i=1}^{d}\Pr\left(\{w^{0};h_{i}^{T}w^{0}=0\}\right)$ (13)
$=$ $0$ (14)
Therefore, $\nabla\mathcal{L}(w^{0})$ is elementwise non-zero w.p.1.., so $C_{l,1}$ and $C_{l,2}$ are non-zero for all $l\in[L]$, w.p.1.. In the following analysis, We will assume $C_{l,1}$ and $C_{l,2}$ are non-zero.
Without loss of generality, we assume $h=0$. This is because minimizing $\mathcal{L}(w)=\frac{1}{2}w^{T}Hw-h^{T}w$ is equivalent to minimizing $\mathcal{L}(w)=\frac{1}{2}\left(w-w^{*}\right)^{T}H\left(w-w^{*}\right)$ where $w^{*}=H^{-1}h$. By a linear transformation $z=w-w^{*}$, Adam for minimizing $\frac{1}{2}\left(w-w^{*}\right)^{T}H\left(w-w^{*}\right)$ starting from $w^{0}$ is equivalent to Adam for minimizing $\frac{1}{2}z^{T}Hz$ starting from $z^{0}=w^{0}-w^{*}$. Thus we can assume $w^{*}=0$, or equivalently, $h=0$. The update rule of Adam becomes
@@
where $D_{Adam}^{0}=\operatorname{diag}(\nabla\mathcal{L}(w^{0})\circ\nabla\mathcal{L}(w^{0}))^{\frac{1}{2}}=\operatorname{diag}(|Hw^{0}|)$. We denote $d_{t}=\eta(D_{Adam}^{0})^{-1}Hw^{t}$ and thus we have $w^{t}=\frac{1}{\eta}H^{-1}D_{Adam}^{0}d^{t}$ and $w^{t+1}=w^{t}-d_{t}$. These relations also hold for each block by changing the notation to $H_{l}$ $w_{l}^{t}$, $D_{Adam}^{0}$, and $d_{l}^{t},$ etc.. Following the framework in [81], we try to bound the error yet to be optimized (a.k.a., cost-to-go) and the per-step improvement, respectively. The ratio of these two terms characterizes the rate of convergence. We now express both terms using $d_{l}^{t}$. For the cost-to-go term for the $l$-th block, we have
@@
$\leq$ $\frac{\max_{l\in[L]}C_{l,2}^{2}}{\min_{l\in[L]}C_{l,1}^{2}}\kappa_{l},$ (19)
@@
$\leq$ $\sum_{l=1}^{L}\left(1-\frac{1}{\left(\frac{\max_{l\in[L]}C_{l,2}^{2}}{\min_{l\in[L]}C_{l,1}^{2}}\right)\kappa_{l}}\right)\left([\mathcal{L}(w^{t})]_{l}-[\mathcal{L}^{*}]_{l}\right)$
$\leq$ $\max_{l\in[L]}\left(1-\frac{1}{\left(\frac{\max_{l\in[L]}C_{l,2}^{2}}{\min_{l\in[L]}C_{l,1}^{2}}\right)\kappa_{l}}\right)\sum_{l=1}^{L}\left([\mathcal{L}(w^{t})]_{l}-[\mathcal{L}^{*}]_{l}\right)$
$=$ $\max_{l\in[L]}\left(1-\frac{1}{\left(\frac{\max_{l\in[L]}C_{l,2}^{2}}{\min_{l\in[L]}C_{l,1}^{2}}\right)\kappa_{l}}\right)\left(\mathcal{L}(w^{t})-\mathcal{L}^{*}\right).$